OpenStudy (anonymous):

The sum of two numbers is 52. The greater number is 4 more than the smaller number. Which equation can be used to solve for the smaller number? x − (x + 4) = 52 x + (x + 4) = 52 x(x + 4) = 52 x(x − 4) = 52

OpenStudy (anonymous):

@Killer_Shadow

OpenStudy (ghostgate):

Simply put, 'x' is the smaller number, now you have the 'sum' which leads to addition of two numbers, since the larger number is 4 "more than" the smaller number, you would then have 'x + 4', concluding this, which equation do you think would be the answer?

OpenStudy (anonymous):

i dont know thats why im asking theres 3 x+4 s i dont know

OpenStudy (ghostgate):

Alright, let me narrow it down: x - (x + 4) = 52 - In this equation you have your two numbers that I mentioned earlier, 'x' which is the smaller number and '(x + 4)' which is the larger number, now the 'sum of two numbers' references the 'small number' and the 'large number' which is 'x' and '(x + 4)', with that in mind now as I said the 'sum' leads to addition, not subtraction. x(x + 4) = 52 and x(x - 4) = 52 - Both deal with multiplication which would conclude the 'product' not the 'sum' of two numbers, so this leaves you with...: x + (x + 4) = 52 - Why? Because your smaller number 'x' is added to the larger number '(x + 4) which then equals the sum of two numbers which would be '52'. Does that make more sense? {---Ghostgate---}

OpenStudy (ghostgate):

In conclusion the answer is: x + (x + 4) = 52 x = 24

OpenStudy (anonymous):

that helped thx

OpenStudy (ghostgate):

Well I'm glad it did, have a great day! {---Ghostgate---}